OpenStudy (anonymous):

how to find that a function f(x) is continuous in [a,b] and differentiable in (a,b)

OpenStudy (amistre64):

i believe that in order for f(x) to be differentiable then the \(\lim_{x->c}\) f(c) from the left and right have to match.

OpenStudy (amistre64):

also, for continuity, check for aymptonic behaviours perhaps

OpenStudy (anonymous):

it must fulfill 3 conditions for it to be continuous for f(x) at point x=c. f(c) must exist, limit of f(x) as x approaches c must also exist. and the limit of left and right must be the same. For it to be differentiable, the function must also be continuous, but if the function has a sharp corner, you can draw more than 1 tangent line at that point, and because the slopes of their tangent lines are not equal, the function will not be differentiable.

OpenStudy (anonymous):

Example: Show the following function is continuous, but NOT differentiable. (the idea is the same) \[f(x)=2x+3, (x <1) \] \[f(x)=3x+2, (x \ge1) \] we see that \[\lim_{x \rightarrow 1^{-}} f(x)= \lim_{x \rightarrow 1^{-}} (2x+3)=2+3=5=3+2=\lim_{x \rightarrow 1^{+}} (3x+2)=\lim_{x \rightarrow 1^{+}} f(x) \] This implies that \[\lim_{x \rightarrow 1} f(x)=5. \] we also see f(1)=3+2=5. Now we have \[\lim_{x \rightarrow 1} f(x)=f(1)\] which proves continuity at 1. The only point where continuity could exist possibly in doubt. Now we show that f'(1) does not exist. if f'(1) existed , then the derivative of 2x+3 = 2 and derivative of 3x+2 = 3 would imply that \[2\neq3\], there are 2 different slopes, thus the function is not differentiable.